# Robust-Wide: Robust Watermarking against Instruction-driven Image Editing

### v1

@@ 4.4 More Analysis
Figure 7: The influence of the number of sampling steps. (a) Some visual examples at different steps and their corresponding BER. (b) BER decreases with more sampling steps.
Figure 8: Scatter plots to show the relationship between the image editing strength and watermark extraction rate.
Relationship between editing strength and extraction ability. We explore the impact of the image editing strength (quantified by PSNR and SSIM computed between $I_{wm}$ and $I^{edit}_{wm}$) on the extraction BER. As shown in Figure 8, points with higher BER are mainly at the area where PSNR and SSIM are low, i.e., the editing strength is large. This is intuitive as the greater the change is, the more difficult it will be for the watermark extraction. When the editing strength on the original image is significant, the resulting edited image can be considered as a form of re-creation. In such cases, the risk of copyright infringement is relatively low, and the ineffective extraction may be deemed acceptable.
@@
Embedding mode of Robust-Wide. Figure 9 displays the normalized residual image between the watermarked and original images, using different watermarking methods. We can see that the watermarks embedded using DWT-DCT and DWT-DCT-SVD are quite faint and imperceptible, indicating their limited robustness. MBRS tends to concentrate primarily along the object's edge contours, making it vulnerable to changes in the image background. In contrast, RivaGAN and our Robust-Wide embed watermarks in both the edges and backgrounds, resulting in enhanced robustness. Furthermore, we use Discrete Cosine Transform (DCT) and Discrete Fourier Transform (DFT) to visualize the residual images in the frequency domain. Watermarks embedded by DWT-DCT and DWT-DCT-SVD appear relatively uniform and weak. RivaGAN predominantly places watermarks in the low-frequency information areas, while MBRS distribute watermark information uniformly in the frequency domain. In contrast, Robust-Wide focuses on embedding watermarks in both low and high-frequency information regions, potentially making it more robust against semantically and conceptually related image modifications.
Figure 10: The impact of editing on the normalized residual image.
Extracting mode of Robust-Wide. We use the original image $I_{ori}$ and watermarked image $I_{wm}$ to generate the corresponding edited images $I_{ori}^{edit}$ and $I_{wm}^{edit}$. Here, we control all potential random factors (e.g., generation random seed) to guarantee that the difference between $I_{ori}^{edit}$ and $I_{wm}^{edit}$ is solely caused by instruction-driven image editing. As shown in Figure 10, the watermark embedded into the original image also exists after editing, which preserves as an outline of the character. We hypothesize that the extractor mainly focuses on such robust concept-aware areas.
@@ 4.5 Ablation Study
$\lambda_{1}$ BER(%)normal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$
0 1.7391 39.6859 0.9750
0.001 2.6579 41.9142 0.9910
0.01 6.0690 41.7604 0.9938
0.1 50.0364 56.9096 0.9986
Table 5: The impact of different $\lambda_{1}$ values. The gray cell denotes the default setting (same for the following tables).
$\lambda_{2}$ BER(%)normal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$
1 0.9674 35.2577 0.9616
0.1 2.6579 41.9142 0.9910
0.01 11.1796 45.8701 0.9929
Table 6: The impact of different $\lambda_{2}$.
Impact of hyper-parameters $\lambda_{1}$ and $\lambda_{2}$. Table 5 shows the watermark performance with different $\lambda_{1}$ values. We observe that a larger $\lambda_{1}$ can help improve the visual quality of watermarked images but causes a decrease in the watermark extraction rate. Table 6 shows the results with different $\lambda_{2}$ values. A higher $\lambda_{2}$ leads to lower BER, which showcases the trade-off between extraction ability and visual quality.
$L_{ex_{2}}$ BER(%)normal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$
w/o Editing w/ Editing
w/o. 50.1098 50.1219 68.7220 0.9999
w/. 0.0000 2.6579 41.9142 0.9910
Table 7: The importance of $L_{ex_{2}}$.
@@
# of Gradient Backward Steps BER(%)normal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$
1 4.0520 44.3330 0.9938
2 2.9166 42.1386 0.9919
3 2.6579 41.9142 0.9910
Table 8: The influence of the number of gradient backward steps.
Influence of the number of gradient backward steps $k$. Table 8 shows the watermark performance with different numbers of gradient backward steps. With more steps, the watermark message is easier to extract while the visual quality of the watermarked image slightly reduces. Due to the GPU memory constraints, the maximum number of gradient backward steps we could set is 3, which is sufficient to obtain acceptable performance.
Watermark lengths (bits) BER(%)normal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$
16 2.2812 40.8327 0.9853
64 2.6579 41.9142 0.9910
256 4.1867 39.1842 0.9844
1024 6.3036 36.5956 0.9732
Table 9: The influence of different watermark message lengths.
Influence of different watermark lengths. Table 9 reports the evaluation results with different lengths of watermark messages. A longer watermark message results in higher BER and lower visual quality. In practice, the user can customarily select the watermark lengths to balance such trade-off. In Table 9, we can observe that the visual quality is lower when the watermark length is 16 compared to when the watermark length is 64. Indeed, the integration of watermark bit with the image involves shape transformations through convolutional or deconvolutional layers. With 16 bits, our available GPU memory posed restrictions on employing convolutional or deconvolutional layers for shape transformations. Consequently, we opted for a non-parametric interpolation method to handle the integration, causing degradation in visual quality.
@@ 5 Conclusion
In this paper, we propose Robust-Wide, the first robust watermarking methodology against instruction-driven image editing. Our core idea is to integrate a novel Partial Instruction-driven Denoising Sampling Guidance (PIDSG) module into the encoder-decoder training framework. Experiments demonstrate that Robust-Wide is robust against instruction-driven image editing in different scenarios while maintaining high visual quality and editability. Our in-depth analysis on the embedding and extracting modes of Robust-Wide is expected to shed light on the design of watermarking against other semantic distortions.
Appendix A Real-World Images Dataset
Here, we present a comprehensive overview of the construction process for the real-world images dataset. To encompass a diverse range of image modifications reflecting real-world scenarios, we opted for six categories of images from the Internet: person, animal, object, architecture, painting, and scenery. For each type, we chose five images and designed six hand-crafted editing instructions, respectively. More details are provided in the following part.
A.1 Real-World Images of Different Types
We show real-world images of different types in Figure 11.
Figure 11: The real-world images of six types.
A.2 Hand-crafted Editing Instructions
We show the designed editing instructions for each type.
Animal
``Move it to the desert"
``Put on a pair of sunglasses"
``Make it a cartoon"
``Make it in Van Gogh's artistic style"
``Make it Picasso style"
``Make it Minecraft"
Architecture
``Make it an Egyptian sculpture"
``Turn it into ruins"
``Put it on fire"
``Make it Picasso style"
``Make it in Van Gogh's artistic style"
``Make it Minecraft"
Object
``Replace it with an apple"
``Change it with a rose"
``Move it on the moon"
``Make it a Modigliani painting"
``Turn this into the space age"
``Make it Minecraft"
Painting
``Make it a Modigliani painting"
``Make it a Miro painting"
``Make it Picasso style"
``Make it a Van Gogh's painting"
``Make it Minecraft"
``Make it a cartoon"
Person
``Apply face paint"
``Make it a zombie"
``Take off the clothes"
``Make it a marble roman sculpture"
``It should look 100 years old"
``Make it Minecraft"
Scenery
``Add a beautiful sunset"
``Cover it with snow"
``Make it a cyberpunk painting"
``Make it a Van Gogh's painting"
``Make it Minecraft"
``Make it a pencil drawing"
Appendix B More Visual Examples
B.1 Main Results.
See Figure 12.
Figure 12: More visual examples from real-world images of six types, namely, animal, architecture, object, painting, person, and scenery. For each, the first row shows Original, Normalized Residual, and Watermarked Images and the second row shows Edited Images with the corresponding instructions, while BER is noted below.
B.2 Robustness Against Different Instruction-driven Image Editing Methods.
See Figure 13.
Figure 13: More visual examples of our robustness against different instruction-driven image editing methods.
B.3 Robust-Wide's Extracting Mode.
See Figure 14.
Figure 14: More visual examples of Robust-Wide's extracting mode.
Appendix C Code
Our code will be released upon acceptance.

### v3

@@ 4.4 More Analysis
Embedding mode of Robust-Wide. Figure 6 displays the normalized residual image between the watermarked and original images, using different watermarking methods. We can see that the watermarks embedded using DWT-DCT and DWT-DCT-SVD are quite faint and imperceptible, indicating their limited robustness. MBRS tends to concentrate primarily along the object’s edge contours, making it vulnerable to changes in the image background. In contrast, RivaGAN and our Robust-Wide embed watermarks in both the edges and backgrounds, resulting in enhanced robustness. Furthermore, we use Discrete Cosine Transform (DCT) and Discrete Fourier Transform (DFT) to visualize the residual images in the frequency domain. From the residual images, we can observe that DWT-DCT and DWT-DCT-SVD introduce input-agnostic modification while other methods introduce input-aware modification. From the last two rows of Figure 6, we find that DWT-DCT, DWT-DCT-SVD, and RivaGAN tend to mainly modify the low-frequency area while MBRS, CIN, PIMoG, and SpeMark prefer to embed the watermark into both low-frequency and high-frequency areas. Differently, Robust-Wide mainly focuses on embedding watermarks in infra-low frequency areas or mid-low frequency areas, potentially making it more robust against semantically and conceptually related image modifications.
Figure 7: Example of our extracting mode.
Figure 8: The influence of the number of sampling steps. (a) Some visual examples at different steps and their corresponding BER. (b) BER decreases with more steps.
Extracting mode of Robust-Wide. We use the original image $I_{ori}$ and watermarked image $I_{wm}$ to generate the corresponding edited images $I_{ori}^{edit}$ and $I_{wm}^{edit}$. Here, we control all potential random factors (e.g., generation random seed) to guarantee that the difference between $I_{ori}^{edit}$ and $I_{wm}^{edit}$ is solely caused by instruction-driven image editing. As shown in Figure 7, the watermark embedded into the original image also exists after editing, which preserves as an outline of the character. We hypothesize that the extractor mainly focuses on such robust concept-aware areas. More examples are provided in the supplementary material.
@@
Relationship between editing strength and extraction ability.
Figure 9: The relationship between image editing strength and extraction ability.
We explore the impact of the image editing strength (quantified by PSNR and SSIM computed between $I_{wm}$ and $I^{edit}_{wm}$) on the extraction BER. As shown in Figure 9, points with higher BER are mainly at the area where PSNR and SSIM are low, i.e., the editing strength is large. This is intuitive as the greater the change is, the more difficult it will be for the watermark extraction. When the editing strength on the original image is significant, the resulting edited image can be considered as a form of re-creation. In such cases, the risk of copyright infringement is relatively low, and the ineffective extraction may be deemed acceptable.
@@ 4.5 Ablation Study
Table 5: The importance of $L_{ex_{2}}$. The gray cell denotes the default setting.
Metrics w/o $L_{ex_{2}}$ w/ $L_{ex_{2}}$
BER(%$\downarrow$ w/o editing 50.1098 0.0000
w/ editing 50.1219 2.6579
[…]
Impact of hyper-parameters $\lambda_{1}$ and $\lambda_{2}$. Table 4.5 shows the watermark performance with different $\lambda_{1}$ and $\lambda_{1}$ values. We observe that a larger $\lambda_{1}$ can help improve the visual quality of watermarked images but causes a decrease in the watermark extraction rate. On the other hand, a higher $\lambda_{2}$ leads to lower BER, which showcases the trade-off between extraction ability and visual quality.
Table 6: The impact of different $\lambda_{1}$ and $\lambda_{2}$.
Metrics $\lambda_{1}$ $\lambda_{2}$
0 0.001 0.01 0.1 0.01 0.1 1
BER(%$\downarrow$ 1.7391 2.6579 6.0690 50.0364 11.1796 2.6579 0.9674
[…]
Table 7: The influence of $k$ and the bits length.
Metrics $k$ Bits Length
1 2 3 16 64 256 1024
BER(%$\downarrow$ 4.0520 2.9166 2.6579 2.2812 2.6579 4.1867 6.3036
[…]
Influence of the number of gradient backward steps $k$. Table 4.5 shows the watermark performance with different numbers of gradient backward steps k. With more steps, the watermark message is easier to extract while the visual quality of the watermarked image slightly reduces. Due to the GPU memory constraints, the maximum number of gradient backward steps we could set is 3, which is sufficient to obtain acceptable performance.
Influence of different watermark bits length. Table 4.5 reports the evaluation results with different lengths of watermark messages. A longer watermark message results in higher BER and lower visual quality. In practice, the user can customarily select the watermark lengths to balance such trade-off. In Table 4.5, we can observe that the visual quality is lower when the watermark length is 16 compared to when the watermark length is 64. Indeed, the integration of watermark bit with the image involves shape transformations through convolutional or deconvolutional layers. With 16 bits, our available GPU memory posed restrictions on employing convolutional or deconvolutional layers for shape transformations. Consequently, we opted for a non-parametric interpolation method to handle the integration, causing degradation in visual quality.
@@ 5 Conclusion
In this paper, we propose Robust-Wide, the first robust watermarking method against instruction-driven image editing. Our core idea is to integrate a novel module called PIDSG into the encoder-noise layer-decoder training framework, which forces watermarks embedded in the semantic level. Experiments demonstrate that Robust-Wide is robust against different image editing methods while maintaining high visual quality and editability. Furthermore, our in-depth analysis on the embedding and extracting modes of Robust-Wide is expected to shed light on the design of watermarking against other semantic distortions.
Acknowledgements
This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s), and Singapore Ministry of Education (MOE) AcRF Tier 2 MOE-T2EP20121-0006.